OpenStudy (anonymous):

how can I determine if this limit is convergent or divergent, and if convergent give the limit: lim(n->inf)=(2^n)/(1+n^2)

OpenStudy (anonymous):

try Hospital rule

OpenStudy (anonymous):

tried it but nothing comes out. any other idea?

OpenStudy (anonymous):

when I use the rule, it's divergent

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty}=\frac{ 2^{n} }{ 1+n^{2} }\]

OpenStudy (anonymous):

=[(2^n)ln2]/2n = [(2^n)(ln2)^2]/2 = infitive

OpenStudy (anonymous):

thank you